This is an Example of a Title with only the Main Words in Capitals

First Author

Institute or Department (optional), University or College City and Country

Second Author

Institute or Department (optional), University or College City and Country

Abstract

The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied.

The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied. The text of the abstract, if there is one, or a 3 line space if an abstract is not supplied.

1 Comments

Your cooperation in complying with the instructions provided in example.tex will add considerably to the overall appearance of the volume. Any suggestions that could improve the admittedly crude WICSBOOK.sty file would be appreciated and should be addressed to the publisher.

2 Example of a second level heading (the title is the first level)

2.1 Example of a third level heading

2.1.1 Example of a fourth level heading

A fifth level heading is indicated by a line space above with the heading in normal type not indented.

A fifth level heading is indicated by a line space above with the heading in normal type not indented. Example of an equation

x = \frac{y + 2/2}{y^{2} + 1}.

The text following an equation is automatically indented unless the following is adopted

x = \frac{y + 2/2}{y^{2} + 1}

in which case the text is flush left. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. Example of a reference. Either of LaTEX’s systems may be adopted [1]. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not

2.2 Example of a third level heading

indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented.

A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented. A fifth level heading is indicated by a line space above with the heading in normal type not indented.

References

[1] A. Author, Paper title, Publication (if book then page numbers, publisher details, date of publication) (if Journal then vol. and page numbers